# New rigidity theorem of Einstein manifolds and curvature operator of the second kind

### v1

@@ Abstract.
Using Bochner techniques, we prove that a compact Einstein manifold of dimension $n\geq 4$ has constant curvature provided that the curvature operator of the second kind satisfies a cone condition that is strictly weaker than nonnegativity. Furthermore, employing a result of Li [Li5], we establish that any closed Einstein manifold of dimension $n\geq 4$ satisfying
@@
must be either flat or a spherical space form. Here, ${\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{\frac{(n-1)(n+2)}{2}}$ are the eigenvalues of $\mathring{R}\,$, $\bar{\lambda}$ is their average, and $\theta(n,k)$ is a positive constant. This result generalizes the work of Dai-Fu [DF] and Chen-Wang [CW1, CW].
@@
The study of the Riemannian curvature operator of the second kind originated with Nishikawa’s conjecture [Ni], which states that a closed simply connected Riemannian manifold with a nonnegative curvature operator of the second kind is diffeomorphic to a locally symmetric space, and if the operator is positive, the manifold is diffeomorphic to a spherical space form. This conjecture was confirmed by Cao–Gursky–Tran [CGT] under the assumption of two-positivity, and later strengthened by Li [Li4] and Nienhaus–Petersen–Wink [NP] under the weaker condition of three-nonnegativity. Further classification results under various curvature conditions of the second kind can be found in [DFD, Li3, Li7, Li6, Li4, Li2, Li5, DF, NP].
@@
Recently, Li [Li5] studied cone conditions of the curvature operator of the second kind, which satisfies
@@
for its eigenvalue of curvature operator of the second kind. This condition is denote by $C(\alpha,\theta)$ and $\mathring{C}\,(\alpha,\theta)$ if the inequality is strict. Li proved that if $\mathring{R}\,\in C(\frac{n+2}{2},\theta)$ for some $1<\theta<\frac{2(n-1)}{n+2},$ then $M$ is either flat or a rational homology sphere.
Building on the work of Li [Li5] and Dai–Fu [DF], Cheng and Wang [CW] relaxed the curvature assumption to a cone condition. They introduced a positive constant $\theta(n)$ and proved that if an $n$-dimensional closed Einstein manifold ($n=4,5$, or $n\geq 8$) satisfies $\lambda_{1}+\lambda_{2}\geq-2\theta(n)\bar{\lambda}$, then it is either flat or a spherical space form. This condition is strictly weaker than $2$-nonnegativity, which corresponds to $\theta=0$.
In this paper, we investigate Einstein manifolds of dimension $n\geq 4$ whose curvature operator of the second kind satisfies a cone condition that is strictly weaker than nonnegativity. Let $\{\lambda_{i}\}_{i=1}^{N}$ be the eigenvalues of $\mathring{R}\,$ and $\bar{\lambda}=\sum\nolimits_{i=1}^{N}{{\lambda}_{i}}/{N}\;$ their average. By applying results of Dai–Fu [DF] and Li [Li5], we obtain the following theorems.
@@
Let $(M^{n},g)$ be an $n$-dimensional compact Einstein manifold. For $n\geq 8,n\neq 10$ and $k\leq[\frac{n+2}{4}]$, there exists a constant $\theta(n,k)>0$ such that if the curvature operator of the second kind satisfies
@@
When $n=4,5$ we have the following theorem.
@@
Let $(M,g)$ be an $n=4$ or $5$-dimensional compact Einstein manifold, and $\mathring{R}\,$ be the curvature operator of the second kind. If
(i) $n=4$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{17}{7}\bar{\lambda}$ or ${\lambda}_{1}+{\lambda}_{2}\geq-\bar{\lambda}$, or
(ii) $n=5$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{47}{18}\bar{\lambda}$ or ${\lambda}_{1}+{\lambda}_{2}\geq-\frac{4}{3}\bar{\lambda}$ or ${\lambda}_{1}+{\lambda}_{2}+{\lambda}_{3}\geq-\frac{1}{18}\bar{\lambda}$,
@@
For dimensions $n=6,7$, we have $\frac{n+2}{4}]=2$, yet our method does not yield a corresponding cone condition for $k=2$. Similarly, for $n=10$, the method does not provide a cone condition for $k=3$. Nevertheless, the following theorem remains valid for the cases $n=6,7$ with $k=1$, and for $n=10$ with $k=1,2$.
@@
Let $(M,g)$ be an $n=6,7$ or $10$-dimensional compact Einstein manifold, and $\mathring{R}$ be the curvature operator of the second kind. If
@@
When $\theta=0$, Theorem A,B,C are the result of Dai-Fu [DF]. When $k=1,2$, Theorem A,B,C are the result of Chen-Wang [CW1, CW].
2. notation and lemmas
Here, we introduce our notation and summarize key facts regarding the curvature operator of the second kind. Further details can be found in [Li4, Li5, NP, DF, CW].
@@
$\bar{R}$ induces a symmetric form $\mathring{R}\,$ defined by
@@
which is called curvature operator of the second kind. It was introduced by Nishikawa [Ni] and as pointed out in [NP], the curvature operator of the second kind $\mathring{R}\,$ can also be interpreted as $\bar{R}$ and its image restrict on $S_{0}^{2}(V)$.
Let $(M,g)$ be an $n$-dimensional Riemannian manifold. For each $p\in M,$ let $V=T_{p}M$ and $R$ is denoted by the Riemannian curvature tensor of $M$. $\mathring{R}\,$ is denoted by curvature operator of second kind of $M$. We introduce the following definition which comes from [NP]
@@
$\sum\limits_{i=1}^{N}{{\lambda}_{i}\left|S^{i}W\right|^{2}}$ $\geq$ $\sum\limits_{i=1}^{k}{{\lambda}_{i}\left|S^{i}W\right|^{2}}+\sum\limits_{i=k+1}^{N}{{\lambda}_{k+1}\left|S^{i}W\right|^{2}}$ (2.1)
@@
${{\lambda}_{k+1}}\geq\frac{{\lambda}_{1}+\cdots+{\lambda}_{k}}{k}\geq-\theta\bar{\lambda}.$ (2.2)
@@
3. proof of main theorem
Let $(M^{n},g)$ be an Einstein manifold of dimension $n$. Denote by $\{\lambda_{i}\}_{i=1}^{N}$ the eigenvalues of the operator $\mathring{R}\,$ and let $\bar{\lambda}=\sum\nolimits_{i=1}^{N}{{\lambda}_{i}}/{N}\;$ be their average. From [Li4], it follows that
@@
Based on the calculations presented in [DF], it has been calculated that
$S=\sum\limits_{i=1}^{N}{{{\left|{{S}^{i}}W\right|}^{2}}}=\frac{2({{n}^{2}}+n-8)}{n}{{\left|W\right|}^{2}}=\frac{4N-12}{n}{{\left|W\right|}^{2}},$ (3.1)
@@
${{\left|W\right|}^{2}}=\frac{4}{3}\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}-\frac{4N}{3}{{\bar{\lambda}}^{2}}.$ (3.2)
@@
$3\left\langle\Delta R,R\right\rangle$ $\geq$ $16\sum\limits_{i=1}^{N}{\lambda_{i}^{3}}+\frac{16}{3n}[2N-12n+6-(N-3)\theta]\bar{\lambda}\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}$ (3.3)
@@
The following Bochner formula comes from Proposition 3.4 in [DF].
@@
When $k=2$, this is Lemma 3.3 of [CW].
@@
Let $0\leq{\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{N}$ and $\sum\limits_{i=1}^{N}{{\lambda}_{i}}=C>0$. If $C=\frac{N(N-1)}{2N-1},$ then the function
@@
$\mu=\frac{1}{3}[1-{\left(\frac{3C-N}{N-2k}\right)^{2}}].$ (3.5)
@@
$\underset{\partial\Omega}{\mathop{\min}}\,F=\underset{1\leq k<N}{\mathop{\min}}\,\frac{C^{3}}{(N-k)^{2}}-\frac{C^{2}}{N-k}.$
@@
$\underset{\partial\Omega}{\mathop{\min}}\,F=\underset{1\leq k<N}{\mathop{\min}}\,\frac{C^{3}}{(N-k)^{2}}-\frac{C^{2}}{N-k}=\frac{C^{3}}{N^{2}}-\frac{C^{2}}{N},$
@@
If $\bar{\lambda}=0$, then $M$ is flat. Now we consider $\bar{\lambda}>0$.
@@
$(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})=(\bar{\lambda},\bar{\lambda},\ldots\bar{\lambda})$ (3.7)
@@
$(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})=(-D\bar{\lambda},\frac{N+D}{N-1}\bar{\lambda},\ldots,\frac{N+D}{N-1}\bar{\lambda}).$ (3.8)
@@
Therefore, $M$ is a locally symmetric space. Since when $n\geq 8,$ $\theta(n,k)$ defined by (3.6) satisfies $\theta<\frac{2(n-1)}{n+2}$ for $1\leq k\leq\left[\frac{n+2}{4}\right]$. It follows from Theorem 1.8 and Proposition 2.9 of [Li5] that $M$ is either flat (when $\bar{\lambda}=0$) or a rational homology sphere (when $\bar{\lambda}>0$).
The complete classification of compact symmetric spaces that are rational homology spheres is due to Wolf [Wo, Theorem 1]. Other than spheres, the only simply-connected example is the space SU(3)/SO(3).{SU(3)}/{SO(3)}. (3.7) and (3.8) implies that the eigenvalues of the curvature operator are either
@@
where $D=\frac{(N-1)k\theta+N(k-1)}{N-k}$. In conjunction with the result of [NP, Example 4.5], it follows that the manifold must be a spherical space form, which completes the proof of Theorem A
@@
When $n=4,5$, according to [DF], due to Jack and Parker’s result [JP] we have
$\left\langle\Delta R,R\right\rangle=8\sum\limits_{i}{\lambda_{i}^{3}}+\frac{8(n-4)}{3}\bar{\lambda}\sum\limits_{i}{\lambda_{i}^{2}}-\frac{4(n+2)(n-1)^{2}}{3}{\bar{\lambda}^{3}}.$
[…]
$\left(-\frac{17}{7}\bar{\lambda},\frac{10}{7}\bar{\lambda},\dots,\frac{10}{7}\bar{\lambda}\right)\quad\text{if}\quad n=4,$
and
$\left(-\frac{47}{18}\bar{\lambda},\frac{23}{18}\bar{\lambda},\dots,\frac{23}{18}\bar{\lambda}\right)\quad\text{if}\quad n=5.$
It follows from [NP, Theorem B] that $M$ is either flat (when $\bar{\lambda}=0$) or a rational homology sphere (when $\bar{\lambda}>0$). The remainder of the argument is analogous to that in Theorem A, which completes the proof of Theorem B. ∎

### v2

@@ Abstract.
Using Bochner techniques, we prove that an Einstein manifold of dimension $n\geq 4$ has constant curvature provided that the curvature operator of the second kind satisfies a cone condition that is strictly weaker than nonnegativity. Furthermore, employing a result of Li [Li25], we establish that any closed Einstein manifold of dimension $n\geq 4$ satisfying
@@
must be either flat or a spherical space form. Here, ${\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{\frac{(n-1)(n+2)}{2}}$ are the eigenvalues of $\mathring{R}\,$, $\bar{\lambda}$ is their average, and $\theta(n,k)$ is a positive constant. This result generalizes the work of Dai-Fu [DF24] and Chen-Wang [CW25, CW24]. We also classify four-dimensional Einstein manifolds satisfying a cone condition.
@@
The study of the Riemannian curvature operator of the second kind originated with Nishikawa’s conjecture [Nis86], which states that a closed simply connected Riemannian manifold with a nonnegative curvature operator of the second kind is diffeomorphic to a locally symmetric space, and if the operator is positive, the manifold is diffeomorphic to a spherical space form. This conjecture was confirmed by Cao–Gursky–Tran [CGT23] under the assumption of two-positivity, and later strengthened by Li [Li24a] and Nienhaus–Petersen–Wink [NPW23] under the weaker condition of three-nonnegativity. Further classification results under various curvature conditions of the second kind can be found in [DFY24, Li22, Li23a, Li23b, Li24a, Li24b, Li25, DF24, NPW23].
@@
Recently, Li [Li25] studied cone conditions of the curvature operator of the second kind, which satisfies
@@
for the first $\alpha$ smallest eigenvalues of the curvature operator of the second kind. This condition is denote by $C(\alpha,\theta)$ and $\mathring{C}\,(\alpha,\theta)$ if the inequality is strict. Li proved that if $\mathring{R}\,\in C(\frac{n+2}{2},\theta)$ for some $1<\theta<\frac{2(n-1)}{n+2},$ then $M$ is either flat or a rational homology sphere.
Building on the work of Li [Li25] and Dai–Fu [DF24], Cheng and Wang [CW24] relaxed the curvature assumption to a cone condition. They introduced a positive constant $\theta(n)$ and proved that if an $n(\geq 8)$-dimensional closed Einstein manifold satisfies $\lambda_{1}+\lambda_{2}\geq-2\theta(n)\bar{\lambda}$, then it is either flat or a spherical space form. This condition is strictly weaker than $2$-nonnegativity, which corresponds to $\theta=0$.
In this paper, we investigate Einstein manifolds of dimension $n\geq 4$ whose curvature operator of the second kind satisfies a cone condition that is strictly weaker than nonnegativity. Let $\{\lambda_{i}\}_{i=1}^{N}$ be the eigenvalues of $\mathring{R}\,$ with arranged in non-decreasing order, and $\bar{\lambda}=\sum\nolimits_{i=1}^{N}{{\lambda}_{i}}/{N}\;$ their average. By applying results of Dai–Fu [DF24] and Li [Li25], we obtain the following theorems.
@@
Let $(M^{n},g)$ be an $n$-dimensional Einstein manifold. For $n\geq 4,n\neq 6,7,10$ and $k\leq[\frac{n+2}{4}]$, there exists a constant $\theta(n,k)>0$ such that if the curvature operator of the second kind satisfies
@@
Theorem  A admits an enhancement for dimensions $n=4$ and $5$, as stated below.
@@
Let $(M,g)$ be an $n=4,5$-dimensional Einstein manifold, and $\mathring{R}$ be the curvature operator of the second kind. If
(i) $n=4$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{1}{7}\bar{\lambda}$,
(ii) $n=5$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{4}{9}\bar{\lambda}$,
@@
For the case $n=4$, Li established a more general cone condition covering all $1\leq k<9$; see [Li25, Throrem 1.5]. Although this condition is stronger than that of Li, we handle it using the Bochner technique. For completeness, we include the details here. As will be seen from the subsequent Theorem D and Corollary 4.1, our approach achieves the optimal result.
Remark 1.2.
For dimensions $n=6,7$, we have $\frac{n+2}{4}]=2$, but our method does not yield a corresponding cone condition for $k=2$. Similarly, for $n=10$, the method does not provide a cone condition for $k=3$. Nevertheless, the following theorem remains valid for $n=6,7$ with $k=1$, and for $n=10$ with $k=1,2$.
@@
Let $(M,g)$ be an $n=6,7$ or $10$-dimensional Einstein manifold, and $\mathring{R}$ be the curvature operator of the second kind. If
@@
When $\theta=0$, Theorem A,B,C are the result of Dai-Fu [DF24]. When $k=1,2$, Theorem A,B,C are the result of Chen-Wang [CW25, CW24]. We note that a sign error in [DF24, (4.9)] led to an incorrect Bochner formula for dimensions $4$ and $5$. This has now been corrected in (3). Hence when $n=4,5$, we can only choose $k=1$.
Based on Lemma $3.3$, we present a rigidity theorem for Einstein $4$-manifolds in the case $k=3$, which improves [Li25, Theorem 1.8] for $n=4$ and establishes its optimality in this case.
Theorem D.
Let $(M,g)$ be an Einstein four-manifold whose the curvature operator of the second kind satisfies
(1.1) $\lambda_{1}+\lambda_{2}+\lambda_{3}\geq-3\bar{\lambda},$
then one of the following must be true:
(1) $M$ is flat.
(2) up to rescaling, $M$ is isometric to a quotient of the round sphere $(\mathbb{S}^{4},g_{0})$.
(3) up to rescaling, $M$ is isometric to a $\mathbb{CP}^{2}$ with the Fubini–Study metric.
(4) up to rescaling, $M$ is isometric to a quotient of $\mathbb{S}^{2}\times\mathbb{S}^{2}$ with the product metric.
2. preliminaries
Here, we introduce our notation and summarize key facts regarding the curvature operator of the second kind. Further details can be found in [Li24a, Li25, NPW23, DF24, CW24].
@@
where $e_{i}\odot e_{j}=e_{i}\otimes e_{i}+e_{j}\otimes e_{i}.$ $\bar{R}$ induces a symmetric form $\mathring{R}\,$ defined by
@@
which is called curvature operator of the second kind. It was introduced by Nishikawa [Nis86] and as pointed out in [NPW23], the curvature operator of the second kind $\mathring{R}\,$ can also be interpreted as $\bar{R}$ and its image restrict on $S_{0}^{2}(V)$.
Let $(M,g)$ be an $n$-dimensional Riemannian manifold. For each $p\in M,$ let $V=T_{p}M$ and $R$ is denoted by the Riemannian curvature tensor of $M$. $\mathring{R}\,$ is denoted by curvature operator of second kind of $M$. We introduce the following definition which comes from [NPW23]
@@
(2.1) $\sum\limits_{i=1}^{N}{{\lambda}_{i}\left|S^{i}W\right|^{2}}$ $\geq$ $\sum\limits_{i=1}^{k}{{\lambda}_{i}\left|S^{i}W\right|^{2}}+\sum\limits_{i=k+1}^{N}{{\lambda}_{k+1}\left|S^{i}W\right|^{2}}$
@@
(2.2) ${{\lambda}_{k+1}}\geq\frac{{\lambda}_{1}+\cdots+{\lambda}_{k}}{k}\geq-\theta\bar{\lambda}.$
@@
3. $n$-dimensional Einstein manifolds
Let $(M^{n},g)$ be an $n$-dimensional Einstein manifold. Denote by $\{\lambda_{i}\}_{i=1}^{N}$ the eigenvalues of the operator $\mathring{R}\,$ and let $\bar{\lambda}=\sum\nolimits_{i=1}^{N}{{\lambda}_{i}}/{N}\;$ be their average. From [Li24a], it follows that
@@ 3. $n$-dimensional Einstein manifolds
Based on the calculations presented in [DF24], it has been calculated that
(3.1) $S=\sum\limits_{i=1}^{N}{{{\left|{{S}^{i}}W\right|}^{2}}}=\frac{2({{n}^{2}}+n-8)}{n}{{\left|W\right|}^{2}}=\frac{4N-12}{n}{{\left|W\right|}^{2}},$
@@
(3.2) ${{\left|W\right|}^{2}}=\frac{4}{3}\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}-\frac{4N}{3}{{\bar{\lambda}}^{2}}.$
@@
(3.3) $3\left\langle\Delta R,R\right\rangle$ $\geq$ $16\sum\limits_{i=1}^{N}{\lambda_{i}^{3}}+\frac{16}{3n}[2N-12n+6-(N-3)\theta]\bar{\lambda}\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}$
@@
The following Bochner formula comes from Proposition 3.4 in [DF24].
@@
When $k=2$, this is Lemma 3.3 of [CW24].
@@
Let $0\leq{\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{N}$ and $\sum\limits_{i=1}^{N}{{\lambda}_{i}}=C>0$, where $N\geq 3$. If $C=\frac{N(N-1)}{2N-1},$ then the function
@@
(3.5) $\mu=\frac{1}{3}[1-{\left(\frac{3C-N}{N-2k}\right)^{2}}].$
@@
$\underset{\partial\Omega}{\mathop{\min}}\,F=\underset{1\leq k<N}{\mathop{\min}}\,\{\frac{C^{3}}{(N-k)^{2}}-\frac{C^{2}}{N-k}\}.$
@@
$\underset{\partial\Omega}{\mathop{\min}}\,F=\underset{1\leq k<N}{\mathop{\min}}\,\{\frac{C^{3}}{(N-k)^{2}}-\frac{C^{2}}{N-k}\}=\frac{C^{3}}{N^{2}}-\frac{C^{2}}{N},$
@@
According to [Li25, Proposition 2.8], we have $\bar{\lambda}\geq 0$ for $\theta(n,k)>0$. If $\bar{\lambda}=0$, then $M$ is flat. Now we consider the case $\bar{\lambda}>0$, i.e., $s>0$. Hence $M$ is compact by Myers Theorem.
@@
(3.7) $(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})=(\bar{\lambda},\bar{\lambda},\ldots\bar{\lambda})$
@@
(3.8) $(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})=(-D\bar{\lambda},\frac{N+D}{N-1}\bar{\lambda},\ldots,\frac{N+D}{N-1}\bar{\lambda}).$
@@
Therefore, $M$ is a locally symmetric space. Since when $n\geq 8,$ $\theta(n,k)$ defined by (3.6) satisfies $\theta<\frac{2(n-1)}{n+2}$ for $1\leq k\leq\left[\frac{n+2}{4}\right]$. It follows from Theorem 1.8 and Proposition 2.9 of [Li25] that $M$ is either flat (when $\bar{\lambda}=0$) or a rational homology sphere (when $\bar{\lambda}>0$).
The complete classification of compact symmetric spaces that are rational homology spheres is due to Wolf [Wol69, Theorem 1]. Other than spheres, the only simply-connected example is the space SU(3)/SO(3).{SU(3)}/{SO(3)}. (3.7) and (3.8) implies that the eigenvalues of the curvature operator are either
@@
where $D=\frac{(N-1)k\theta+N(k-1)}{N-k}$. In conjunction with the result of [NPW23, Example 4.5], it follows that the manifold must be a spherical space form, which completes the proof of Theorem A
@@
The corrected version of [DF24, (4.9)] is as follows.
${R_{tspq}}{R_{tjpl}}{R_{sjql}}=$ ${W_{tspq}}{W_{tjpl}}{W_{sjql}}-\frac{3s}{2n(n-1)}\left|W\right|^{2}+\frac{(n-2)s^{3}}{n^{2}(n-1)^{2}}$
$=$ ${W_{tspq}}{W_{tjpl}}{W_{sjql}}-\frac{3s}{2n(n-1)}\left|R\right|^{2}+\frac{(n+1)s^{3}}{n^{2}(n-1)^{2}}.$
Here the sign for $\frac{{3s}}{{2n(n-1)}}{\left|W\right|^{2}}$ is given as ‘++’ in [DF24, (4.9)], but it should be corrected to ‘-’. Following the approach of [DF24], we present the corrected formula as follows:
$\left\langle\Delta R,R\right\rangle=8\sum\limits_{i}{\lambda_{i}^{3}}+\frac{8(n-10)}{3n(n-1)}s\sum\limits_{i}{\lambda_{i}^{2}}-\frac{4(n-7)(n+2)}{3n^{3}(n-1)^{2}}s^{3}$
(3.9) $=$ $8\sum\limits_{i}{\lambda_{i}^{3}}+\frac{8(n-10)}{3}\bar{\lambda}\sum\limits_{i}{\lambda_{i}^{2}}-\frac{4(n+2)(n-7)(n-1)}{3}\bar{\lambda}^{3}.$
We use the above Bochner formula, and the same methods as the proof of Theorem A, then take $n=4,5$ and $k=1$. We get $n=4$, $\theta=\frac{1}{7}$ and $n=5,\theta=\frac{4}{9}$. The remainder of the argument is analogous to that in Theorem A, which completes the proof of Theorem B. ∎
@@
4. $4$-dimensional Einstein manifolds
In this section, we present a geometric application of Lemma 3.3 in $4$-dimensional Einstein manifolds.
Proof of Theorem D.
The Bochner formula (3) on $4$-dimensional Einstein manifolds can be written as
(4.1) $\left\langle\Delta R,R\right\rangle=8\sum\limits_{\alpha}{\lambda_{\alpha}^{3}}-\frac{4}{3}s\sum\limits_{\alpha}{\lambda_{\alpha}^{2}}+\frac{1}{24}s^{3}.$
Denote the eigenvalues of $W^{+}$ by $a_{1}\leq a_{2}\leq a_{3}$, and those of $W^{-}$ by $b_{1}\leq b_{2}\leq b_{3}$, then
(4.2) $a_{1}+a_{2}+a_{3}=0,b_{1}+b_{2}+b_{3}=0.$
According to [CGT23], on Einstein 4-manifold, the eigenvalues of $\mathring{R}$ are given by
(4.3) ${\lambda}_{ij}=\frac{s}{12}-a_{i}-b_{j},\quad i,j=1,2,3.$
If $\mathring{R}$ satisfies (1.1), then
$a_{3}=\frac{s}{12}-\frac{\lambda_{31}+\lambda_{32}+\lambda_{33}}{3}\leq\frac{s}{12}+\bar{\lambda}=\frac{s}{6},$
and
$b_{3}=\frac{s}{12}-\frac{\lambda_{13}+\lambda_{23}+\lambda_{33}}{3}\leq\frac{s}{12}+\bar{\lambda}=\frac{s}{6}.$
Consequently, $s\geq 0$; moreover, if $s=0$, then $W^{\pm}=0$, i.e., $M$ is flat.
By substituting (4.3) into (4.1) and applying (4.2), we directly compute to obtain
(4.4) $\left\langle\Delta R,R\right\rangle=2[s(a_{1}^{2}+a_{2}^{2}+a_{3}^{2})-12(a_{1}^{3}+a_{2}^{3}+a_{3}^{3})]+2[s(b_{1}^{2}+b_{2}^{2}+b_{3}^{2})-12(b_{1}^{3}+b_{2}^{3}+b_{3}^{3})].$
Let
$f_{1}(a_{1},a_{2},a_{3})=s(a_{1}^{2}+a_{2}^{2}+a_{3}^{2})-12(a_{1}^{3}+a_{2}^{3}+a_{3}^{3}),f_{2}(b_{1},b_{2},b_{3})=s(b_{1}^{2}+b_{2}^{2}+b_{3}^{2})-12(b_{1}^{3}+b_{2}^{3}+b_{3}^{3}).$
Now we consider the case $s>0$. The expression for $f_{1}$ can be rewritten as
$f_{1}=s(a_{1}^{2}+a_{2}^{2}+a_{3}^{2})-12(a_{1}^{3}+a_{2}^{3}+a_{3}^{3})$
$=s\sum\limits_{i=1}^{3}{(\frac{s}{6}-a_{i}-\frac{s}{6})^{2}}+12\sum\limits_{i=1}^{3}{(\frac{s}{6}-a_{i}-\frac{s}{6})^{3}}$
$=12\sum\limits_{i=1}^{3}{(\frac{s}{6}-a_{i})^{3}}-5s\sum\limits_{i=1}^{3}{(\frac{s}{6}-a_{i})^{2}}+\frac{s^{3}}{4}$
$=\frac{125}{144}s^{3}\left[\sum\limits_{i=1}^{3}{[\frac{12}{5s}(\frac{s}{6}-a_{i})]^{3}}-\sum\limits_{i=1}^{3}{[\frac{12}{5s}(\frac{s}{6}-a_{i})]^{2}}\right]+\frac{s^{3}}{4}.$
Since
$\frac{12}{5s}(\frac{s}{6}-a_{i})\geq 0\quad\text{and}\quad\sum\limits_{i}{\frac{12}{5s}(\frac{s}{6}-a_{i})}=\frac{6}{5},$
it follows from Lemma 3.3 that the minimal points of $f_{1}$ are
$(\frac{12}{5s}(\frac{s}{6}-a_{1}),\frac{12}{5s}(\frac{s}{6}-a_{2}),\frac{12}{5s}(\frac{s}{6}-a_{3}))=(\frac{2}{5},\frac{2}{5},\frac{2}{5})$
and
$(\frac{12}{5s}(\frac{s}{6}-a_{1}),\frac{12}{5s}(\frac{s}{6}-a_{2}),\frac{12}{5s}(\frac{s}{6}-a_{3}))=(\frac{3}{5},\frac{3}{5},0),$
which correspond respectively to
$(a_{1},a_{2},a_{3})=(0,0,0)\quad\text{and}\quad(a_{1},a_{2},a_{3})=(-\frac{s}{12},-\frac{s}{12},\frac{s}{6}).$
Hence,
$f_{1}\geq f_{1}(0,0,0)=f_{1}(-\frac{s}{12},-\frac{s}{12},\frac{s}{6})=0.$
Similarly, the minimal points of $f_{2}$ are $(b_{1},b_{2},b_{3})=(0,0,0)$ and $(b_{1},b_{2},b_{3})=(-\frac{s}{12},-\frac{s}{12},\frac{s}{6}),$ and
$f_{2}\geq f_{2}(0,0,0)=f_{2}(-\frac{s}{12},-\frac{s}{12},\frac{s}{6})=0.$
In summary, we have
$\frac{1}{2}\Delta{{\left|R\right|}^{2}}={{\left|\nabla R\right|}^{2}}+\left\langle\Delta R,R\right\rangle\geq 0.$
By the maximum principle, we obtain $\nabla R=0$, and either
[…]
Therefore, the curvature operator $\hat{R}$ of the first kind has eigenvalues $(\frac{s}{12},\frac{s}{12},\frac{s}{12},\frac{s}{12},\frac{s}{12},\frac{s}{12})$, $(\frac{s}{12},\frac{s}{12},\frac{s}{12},0,0,\frac{s}{4})$, $(0,0,\frac{s}{4},\frac{s}{12},\frac{s}{12},\frac{s}{12})$, or $(0,0,\frac{s}{4},0,0,\frac{s}{4})$. Applying the classification of $4$-dimensional symmetric spaces, we complete the proof of Theorem D. ∎
Corollary 4.1.
Let $(M,g)$ be an Einstein four-manifold with $4\frac{1}{2}$-nonnegative curvature operator of the second kind. Then one of the following must be true:
(1) $M$ is flat.
(2) up to rescaling, $M$ is isometric to a quotient of the round sphere $(\mathbb{S}^{4},g_{0})$.
(3) up to rescaling, $M$ is isometric to $\mathbb{CP}^{2}$ with the Fubini–Study metric.
Remark 4.2.
This corollary has already been proven in [CGT23] using a different method.
Proof.
Suppose that ${\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{9}$ are the eigenvalues of $\mathring{R}\,$ and $\bar{\lambda}$ is their average. Then $4\frac{1}{2}$-nonnegative curvature operator of the second kind $\mathring{R}\,$ satisfies
${\lambda}_{1}+{\lambda}_{2}+{\lambda}_{3}\geq-{\lambda}_{4}-\frac{1}{2}{\lambda}_{5}$
$\geq-\frac{\frac{1}{2}{\lambda}_{5}+{\lambda}_{6}+{\lambda}_{7}+{\lambda}_{8}+{\lambda}_{9}}{3}$
$\geq-\frac{{\lambda}_{1}+{\lambda}_{2}+{\lambda}_{3}+{\lambda}_{4}+{\lambda}_{5}+{\lambda}_{6}+{\lambda}_{7}+{\lambda}_{8}+{\lambda}_{9}}{3}$
$=-3\bar{\lambda}.$
Combining Theorem D, we complete the proof of Corollary 4.1 because $\mathbb{S}^{2}\times\mathbb{S}^{2}$ has $4\frac{1}{2}$-negative curvature operator of the second kind. ∎